OpenStudy (jordanloveangel):

What is 3.2 × 10−4 m/h in centimeters per second? A. 8.889 × 10^−6 cm/s B. 8.889 × 10^−5 cm/s C. 1.152 × 10^−5 cm/s D. 1.152 × 10^−7 cm/s

OpenStudy (lasttccasey):

\[3.2 × 10^{−4} \frac{ m }{ h }|\frac{1 h}{3600 s}|\frac{100 cm}{1 m}=8.889×10^{-6}\frac{cm}{s}\]

OpenStudy (anonymous):

\[3.2*10^-4 m/s = 3.2*10^-4 * \frac{ 100cm }{ 3600 }\] \[= 3.2*10^-4 * \frac{ 1 }{ 36 }cm/s \approx 0.8888....*10^-4 cm/s\] \[= 8.889*10^-6 cm/s\] so A

OpenStudy (anonymous):

its 10^-4 and 10^-6

OpenStudy (anonymous):

do u understand?

OpenStudy (jordanloveangel):

yes thx guy